# Prompt-SAW: Leveraging Relation-Aware Graphs for Textual Prompt Compression

### v1

@@ Appendix C Additional Implementation details
Appendix D Experimental Details
D.1 Dataset
@@
statistics GSM8K-AUG NaturalQuestions
No.1 shot No.2 shot No.3 shot No.4 shot
Token number of prompt 422 311 249 242 3040
Number of questions 1319 1319 1319 1319 2654
@@
(i) GSM8k-aug. Originally GSM8K (Cobbe et al., 2021) data set encompasses 8,500 high-quality, linguistically diverse grade school math word problems. We find that existing research poses some limitations on prompt compression experiments GSM8k, i.e., they only allow compressing one COT prompt (1-shot). On the contrary, we design a more comprehensive experimental setting by extending it to $i$-shot setting ($i\in\{1,2,4\}$), i-shot means $i$ chain-of-thought demonstrations in the prompt. However, this setting may result in a huge amount of experiments. To avoid computational overhead, we only choose a small subset of questions in GSM8K for testing (i.e., approximately 30%). In this paper, we call this setting as GSM8k-aug.
(ii) NaturalQuestions. It is a QA dataset that is comprised of real-world queries collected by individuals (Liu et al., 2023). Each question of this dataset has 20 related documents, one of which contains the correct answer. We select documents containing answers as compression targets to examine better the compression performance of different methods on a single document.
D.2 Baseline Models
@@
(iii) GPT4.(Achiam et al., 2023))
We designed five sets of prompts for GPT4 to inspire its ability on prompt compression and reported the best one. Appendix B displays the prompts we employed.
D.3 Evaluation Metrics
@@
Appendix E Additional Results
E.1 Interpretability of Compressed prompts (Examples)
In this section, we report some examples prompts along with prompts compressed by Prompt-SAW and LLMLingua (Jiang et al., 2023a) for a quick comparison in terms of readability and/or interpretability of the compressed prompt. As an example, for the compressed prompt text compressed using LLMLingua in Table E.1, the text ”won twoes forg01 theate women prize:ertMayer” is hard to interpret for humans. On the contrary, the prompt compressed by Prompt-SAW yields comprehensive information units helpful that are not only easy to interpret but are also highly relevant to the question.
@@ E.1 Interpretability of Compressed prompts (Examples)
Document [1](Title: OPEC) Organization of the Petroleum Exporting Countries (OPEC, /oupek/ OH-pek, or OPEP in several other languages) is an intergovernmental organization of 14 nations as of February 2018, founded in 1960 in Baghdad by the first five members (Iran, Iraq, Kuwait, Saudi Arabia, and Venezuela), and headquartered since 1965 in Vienna, Austria. As of 2016, the 14 countries accounted for an estimated 44 percent of global oil production and 73 percent of the world’s p̈rovenöil reserves, giving OPEC a major influence on global oil prices that were previously determined by American-dominated multinational oil companies. Question: how many countries are a part of opec Answer: Compressed Prompt by LLMLingua:
@@
Appendix F Limitations
Although Large Language Models (LLMs) have garnered widespread attention for their remarkable capacity for knowledge comprehension (Yang et al., 2024b; a), enabling tailored solutions across various applications, they also face critical issues such as privacy concerns (Hu et al., 2023a), and explainability (Hu et al., 2023b; Lai et al., 2023). LLM applications often involve data containing sensitive information, necessitating effective solutions to safeguard privacy (Xu et al., 2023a). One promising approach to address this challenge is through the design of Differentially Private (DP) algorithms (Dwork et al., 2006). DP offers provable protection against identification and is resilient to arbitrary auxiliary information that may be available to attackers. While there have been numerous studies on DP machine learning (Hu et al., 2022; Wang et al., 2020; 2021; Su et al., 2022; Hu et al., 2023c; Wang et al., 2023) and DP deep learning (Xiang et al., 2024; 2023; Shen et al., 2023), most of these efforts have primarily focused on either continuous tabular data or image data. Unfortunately, there has been less emphasis on adapting DP algorithms to the context of Natural Language Processing (NLP) and the text domain. Addressing this gap is paramount as text data presents unique challenges and characteristics that necessitate specialized privacy-preserving techniques. By developing and refining DP algorithms tailored to NLP tasks, we can enhance the privacy protections of LLMs and facilitate their responsible and ethical deployment across various domains. However, this endeavor is left for future exploration and research.

### v2

@@ 5 GSM8K-AUG
6 Experiments
6.1 Experiment Settings
Datasets. To comprehensively evaluate the effectiveness of compressed prompts, we evaluate their performance under both task-agnostic and task-aware data settings. For task-agnostic data sets, we consider GSM8K-aug, i.e., an extended variant of the original GSM8K (Cobbe et al., 2021) devised by us to report the model performance under $i$-shot settings with $i\in\{1,2,4,8\}$ (details in Section 5). For the task-aware dataset, we use NaturalQuestions (Liu et al., 2023), and ShareGPT [^2] https://sharegpt.com/. The statistics of dataset is given in Table 5, and further details are provided in Appendix C.1.
Baselines. We compare the performance of Prompt-SAW against following models as baselines: (i) Selective-Context (Li, 2023), (ii) LLMLingua (Jiang et al., 2023a), (iii) LongLLMlingua (Jiang et al., 2023b), and (iv) GPT4 (Achiam et al., 2023). Details about the baselines are provided in Appendix C.2. Note, in order to setup a fair platform for comparative evaluation, we re-compute the results for the baseline models as per our data settings.
Evaluation Metrics. For GSM8K-aug, we use Exact Match (EM) as the evaluation metric. This metric is also employed by Cobbe et al. (2021) and Jiang et al. (2023a). For the evaluation of NaturalQuestions, we used Span Accuracy (Span-Acc) as a metric. This is similar to previous work by Liu et al. (2023) and Jiang et al. (2023b). For the evaluation of ShareGPT, we used Rouge as the evaluation metric (Lin, 2004). Apart from these, we also use fluency (FL) (Meng et al., 2022) to measure the readability and grammatical coherence of the compressed prompt. Further details and mathematical formulation of these metrics are given in Appendix C.3.
Method GSM8K-aug(1-shot) GSM8K-aug(2-shot) GSM8K-aug(4-shot) GSM8K-aug(8shot)
EM Tokens 1/$\eta$ EM Tokens 1/$\eta$ EM Tokens 1/$\eta$ EM Tokens 1/$\eta$
Original 73.33 306 1.00 78.17 612 1.00 78.92 1224 1.00 82.53 2365 1.00
Selective-Context 55.55 228 1.34 58.01 449 1.36 58.13 881 1.39 61.47 1752 1.35
LLMLLingua 63.25(7.3%) 232 1.32 65.70 428 1.43 67.41(5.5%) 906 1.35 73.02(4.2%) 1799 1.31
GPT4-Generation 60.51 215 1.42 66.44(10.1%) 411 1.49 66.39 956 1.28 71.41 1273 1.86
Prompt-SAW 67.84 207 1.48 73.17 399 1.53 71.14 820 1.49 76.13 1586 1.49
Table 1: Experimental results on GSM8K-aug. We report the avg., number of Tokens in original and compressed prompts along with EM and compression ratio $(1/\eta)$. For these results, we use GPT3.5-turbo as target LLM. $\eta^{*}$ is set equal to 0.7. We bold-face overall best scores, and underline the state-of-art along with relative %-age performance improvement.
Large Models. To demonstrate the generalization of our algorithm on different LLMs, we use GPT3.5-turbo and LLaMA2-7B-chat as our target LLMs.
Experimental Setup. Following the setting of LLMLingua (Jiang et al., 2023a), we employ greedy decoding with the temperature set to 0. The max number of tokens generated by LLMs are limited to 400. For graph construction, we use Open-IE tooklit  (Kolluru et al., 2020) as the primary tool and Phi-3-mini  [^3] https://ollama.com/library/phi3 as our auxiliary solution. Note, on average 90-graphs were constructed using the Open-IE toolkit. We use OpenAI embedding API  [^4] https://openai.com/ as the embedding encoder $(E)$. The value for $\eta^{*}$ is set to {0.1, 0.3, 0.5} for both ShareGPT and NaturalQuestions, while $\eta^{*}$ = 0.7 for GSM8K-aug. In Algorithm 2, we use $\gamma$ = 0.001. All the results reported in this paper are averaged over five runs. All experiments were performed using PyTorch 2.1.0 with Nvidia RTX 4090 24GB GPU.
6.2 Experimental Results
Results for Task-agnostic Settings. For task-agnostic settings, we report the results of Prompt-SAW for GSM8K-aug in Table 1. Note, unlike existing research that reports their performance for one fixed setting, we report these results for $i$-shot settings, where $i$ indicates the number of prompts have been employed by Prompt-SAW, i.e., {1, 2, 4 and 8}-shots.
Comparing these results against the baseline models, we can observe that Prompt-SAW outperforms the previous state-of-the-art by a significant margin. For instance, compared to the best performing baselines, Prompt-SAW improves the EM score by up to 7.3%, 10.1%, 5.5% and 4.2% under 1-shot, 2-shot, 4-shot and 8-shot settings, respectively. Correspondingly reduction in the prompt size is 32.3%, 34.9%, 33.0% and 32.9%. We attribute such drastic performance improvement to the following factors: (1) Prompt-SAW retains the logical integrity of the prompts by sub-dividing the original prompts into smaller comprehensive information elements; (2)Prompt-SAW benefits from the workflow that allows selecting and omitting individual information elements for prompt compression without destroying the overall information structure of the compressed prompt. These help Prompt-SAW to ensure the utility of the compressed prompt for the end task.
Target LLM Method NaturalQuestions
$\eta^{*}$ = 0.5 $\eta^{*}$ = 0.3 $\eta^{*}$ = 0.1
Acc Tokens 1/$\eta$ Acc Tokens 1/$\eta$ Acc Tokens 1/$\eta$
GPT3.5-turbo Original 92.18 524 1.00 92.18 524 1.00 92.18 524 1.00
Selective-Context 49.23 283 1.85 45.71 173 3.03 31.12 68 7.69
LongLLMlingua 59.65(39.0%) 270 1.94 52.02(40.8%) 161 3.25 47.14(14.7%) 57 9.21
Prompt-SAW 82.93 227 2.31 73.22 136 3.86 54.07 33 16.08
LLaMA2-7B-chat Original 71.25 524 1.00 71.25 524 1.00 71.25 524 1.00
Selective-Context 40.87 283 1.85 34.26 173 3.03 21.37 68 7.69
LongLLMlingua 43.56(77.1%) 270 1.94 38.87(71.2%) 161 3.25 26.12(72.7%) 57 9.21
Prompt-SAW 77.14 227 2.31 66.53 136 3.86 45.11 33 16.08
Table 2: Experimental results on NaturalQuestions. We bold-face overall best scores, and underline the existing state-of-art along with relative %-age performance improvements.
Target LLM Method ShareGPT
$\eta^{*}$ = 0.5 $\eta^{*}$ = 0.3 $\eta^{*}$ = 0.1
Rouge-1 Rouge-2 Rouge-L Tokens 1/$\eta$ Rouge-1 Rouge-2 Rouge-L Tokens 1/$\eta$ Rouge-1 Rouge-2 Rouge-L Tokens 1/$\eta$
GPT3.5-turbo Selective-Context 36.41 16.48 23.17 312 1.80 34.41 14.81 22.04 217 2.59 33.32 11.87 19.66 95 5.92
LongLLMlingua 38.13(29.3%) 17.07(52.6%) 25.22(29.1%) 305 1.84 36.13(34.9%) 15.61(70.9%) 23.17(38.3%) 202 2.78 34.23(38.6%) 12.64(87.6%) 21.16(32.0%) 90 6.24
Prompt-SAW 49.31 26.04 32.57 176 3.19 48.75 26.68 32.04 146 3.85 47.44 23.71 27.93 61 9.21
LLaMA2-7B-chat Selective-Context 32.24 13.42 23.74 312 1.80 30.78 10.95 22.96 217 2.59 30.15 10.32 19.13 95 5.92
LongLLMlingua 34.51(0.8%) 15.18 27.73 305 1.84 32.97(1.1%) 13.78 26.01 202 2.78 31.26(2.8%) 12.11 23.08 90 6.24
Prompt-SAW 34.81 14.03(8.2%) 24.88(11.4%) 176 3.19 33.34 12.19(13.0%) 24.08(8.0%) 146 3.85 32.15 11.34(6.8%) 20.35(13.4%) 61 9.21
Table 3: Experimental results on ShareGPT. We bold-face overall best scores, and underline the existing state-of-art along with relative %-age performance improvements.
Results for Task-aware Settings. Table 2 reports the performance of Prompt-SAW under task-aware settings on NaturalQuestions using GPT3.5-turbo and LLaMA2-7B-chat as target LLMs. These results show that, Prompt-SAW improves the Span Accuracy by 39.0%, 40.8% and 14.7% for GPT3.5-turbo, and 77.1%, 71.2%, 72.7% for LLaMA2-7B-chat, respectively, for different values of the target compression rates $\eta^{*}=\{0.5,0.3,0.1\}$, against the best-performing baseline (LongLLMLingua). Correspondingly, the reduction in the prompt size is 56.7%, 74.0%, and 93.7% respectively.
The results of Prompt-SAW on ShareGPT (in Table 3) show for GPT3.5-turbo as target LLM, Prompt-SAW improves the Rouge-1 score by up to 29.3%, 34.9% and 38.6%. The performance for Rouge-2 and Rouge-L exhibit a similar behavior. For these results, we also observe for LLaMA2-7B-chat, the improvement in performance is relatively lower compared to that of GPT3.5-turbo. A probable justification in this regard is the fact that LLaMA2-7B-chat is more influenced by the change in context for the compressed prompt. Whereas, higher performance on GPT3.5-turbo indicates that Prompt-SAW preserves the critical information in the prompt.
Correlating the results for both settings, we observe that compared to the task-agnostic scenarios, Prompt-SAW yields better performance for task-aware settings, especially for NaturalQuestions. This is due to the fact that it is more difficult to dig out the latent correlation between information elements within the prompt’s internal structure rather than explicit task-aware correlation extraction.
From Table 2 and Figure 3 we can also find that the performance of Prompt-SAW drops when the value for the target compression rate $(\eta^{*})$ decreases from 0.5 to 0.1. A probable justification that the actual compression ratio of Prompt-SAW is significantly higher than the target compression ratio when the target compression ratio is 10, i.e., $\eta^{*}=0.1$. This is owing to the fact that Prompt-SAW only retains the information elements in $\mathcal{G}$ as the key/basic information units for the compression process. It will delete some entities and relations that may be highly similar to the problem but their overall structure is too long, leading to relatively poor performance for a higher compression ratio.
6.3 Further Analysis
Different Target LLMs. We also analyze the performance of Prompt-SAW using different target LLMs. Comparing the results in Table 2 and Table 3, we can find that GPT3.5-turbo performs better than LLaMA2-7B-chat.
For NaturalQuestions, GPT3.5-turbo achieves up to 19.8% higher Acc scores compared to that of LLaMA2-7B-chat. Likewise for ShareGPT, it achieve up to 47.55% higher value for Rouge-1. We attribute this result to GPT3.5’s stronger ability to understand and comprehend context, resulting in generating higher quality response for the prompts compressed by Prompt-SAW.
Selective-Context LLMLingua Prompt-SAW
FL 5.61 5.74 6.30
Readability of Compressed Prompts. As explained in the introduction (also highlighted in Figure 1), a key limitation of existing prompt compression approaches is the limited readability of the compressed prompt. In order to validate the results of Prompt-SAW in terms of human readability and/or interpretability, we report some example prompts along with prompt compressed using Prompt-SAW and LLMLingua (Jiang et al., 2023a) in Appendix D.1, for a quick comparison. These examples clearly indicate that the prompt compressed by Prompt-SAW exhibit better readability and/or interpretability compared to compressed using LLMLingua.
For instance, as shown in Example D.1 (Appendix D.1), the prompt compressed by LLMLingua encompasses grammatical incoherent text, such as: {“List of Nobelates in The first Prize1 Wilhelmrad, of who received82 in en prize”}. Lack of grammatical coherence significantly undermines the readability and/or interpretability of the compressed prompt, thus impacting its end-utility. Whereas, the prompt compressed by Prompt-SAW, relying on knowledge graph triples exhibits a consistent grammatical sentence structure, as shown in the lower half of Example D.1.
Refer to caption Refer to caption
Figure 3: Different target compress rate between LLMLingua and Prompt-SAW on NaturalQuestions and GSM8K-aug dataset using GPT3.5-turbo as target LLM.
To further support our claims, we also conducted a quantitative comparison. Specifically, we assess the fluency of the compressed prompts through the computation of a weighted mean of bi-gram and tri-gram entropies (Meng et al., 2022). Its computational details are given in Appendix C.3, and result is reported in Table 4. These results show that Prompt-SAW yields relatively higher fluency scores than the baseline models. A lower score for baseline models, e.g., LLMLingua, is attributable to loss of intrinsic semantic relationship between the tokens for the compressed prompt.
Computational Overhead. One of the key objectives of prompt compression is to efficiently reduce the overall computational and corresponding fiscal cost associated with the proprietary LLMs, while at the same time preserving the end-utility of the prompt.
In order to compute the computational overhead of Prompt-SAW, we assume embedding and computing similarity takes a constant amount of time and its cost may be ignored because it is much smaller than that of LLM. Specifically, we use the following formulation to study the computational efficiency of Prompt-SAW:
$c=L\cdot c_{\text{graph}}+(L\cdot\eta^{*})\cdot c_{\text{LLMs}},$ (7)
where $c_{\text{graph}}$ represents the per-token computation load to generate knowledge triples and $c_{\text{LLMs}}$ is per-token computation for target LLMs to generate final output respectively. $L$ represents length of the original prompt, $\eta^{*}$ is the target compress rate, and $c$ represents the total computational overhead for compression.
Following the assumption of Jiang et al. (2023a), we estimate $c_{\text{graph}}$ and $c_{\text{LLMs}}$ based on model parameters, as: $c_{\text{graph}}\approx 0.3/175c_{\text{LLMs}}\approx 0.0017\cdot c_{\text{LLMs}}$ using OpenIE toolkit as the knowledge extraction approach. When $\eta^{*}=0.2$, we have $c\approx 0.2017L\cdot c_{\text{LLMs}}$. This means we can achieve nearly 5x savings in terms of computational resources. For the cases employing a small LLM for graph construction, $c_{\text{graph}}\approx 3.8/175c_{\text{LLMs}}\approx 0.02\cdot c_{\text{LLMs}}$, the computational overhead is $c\approx 0.22\cdot c_{\text{LLMs}}$, which is only slightly higher than the first case.
Overall, these results show the computation efficiency of our model is comparable with that of Jiang et al. (2023a). On the other hand, Prompt-SAW offers much higher benefits, i.e., compressing prompts without distorting their readability for end-users, while at the same time preserving their end-utility to the best possible extent.
7 Conclusion
In this work, we proposed Prompt-SAW that leverages graph structures to infer key information in the prompt in order to come up with a compressed prompt. Experimental evaluation showed that Prompt-SAW outperforms the existing research on by a significant margin. Moreover, Prompt-SAW addressed a key limitation of existing prompt compression approaches, i.e., the compressed prompts are easy to read and understand for end-readers.
8 Limitations
Some of the limitations of the Prompt-SAW  are as follows:
  1. 1.
    Currently our work is focused on compressing prompts that may be reformulated into structured elements, i.e., knowledge graph triplets $(s,r,o)$. We consider its generalization to text segments that may not be organized as graph triplets as a future research direction.
  2. 2.
    The performance of Prompt-SAW relies on the quality of the knowledge graph constructed. We use OpenIE toolkit as our primary graph construction tool. The errors in the graph are propagated in the compressed prompt and may impact the end utility of the compressed prompt.
Ethics Statement
This work fully complies with the ACL Ethics Policy. We declare that there are no ethical issues in this paper, to the best of our knowledge.
Appendix A Background
A.1 Knowledge Graph
A.2 Task-aware Prompts
A.3 Task-agnostic Prompts
A.4 Chain-of-thought Prompt
A.5 Prompt with external documentation
A.6 Token-level Prompt compression
LLMLingua is a token-level prompt compression method that performs compression based on perplexity. It includes a budget controller to calculate the compression ratio of demonstrations and the further compression ratio of each demonstration based on given parameters. LLMLingua uses the LLAMA-2 Touvron et al. (2023) to calculate the perplexity of each token. Finally, LLMLingua compresses the prompt based on the perplexity and compression ratio.
Appendix B Prompts
B.1 Prompts for Graph Construction
Prompts for Graph Construction: Example:
Input:
Output:
B.2 Instructions used for GPT-4 response Generation
The instructions we used in the GPT-4 Generation are shown below:
Appendix C Experimental Details
C.1 Dataset
@@
Statistics GSM8K-aug NaturalQuestions ShareGPT
1shot 2shot 4shot 8shot
Token number of prompt 306 612 1224 2365 3040 562
Number of questions 1319 2654 575
@@
(i) GSM8K-aug. we use GSM8K-aug, an extended version of original GSM8k data set allowing computations for under i-shot settings, i.e., i = {1, 2, 4 and 8}-shots. Its process-flow is explained in Section 5. The statistics of the data set is shown in Table 5.
(ii) NaturalQuestions. It is a QA dataset that is comprised of real-world queries collected by individuals Liu et al. (2023). Each question of this dataset has 20 related documents, one of which contains the correct answer. We select documents containing answers as compression targets to examine better the compression performance of different methods on a single document.
(iii) ShareGPT. It is a conversation dataset encompassing users’ conversation with ChatGPT [^5] ShareGPT.com. Each data sample is a complete conversation between the user and ChatGPT, covering multiple languages across different scenarios. Following Li (2023) and Jiang et al. (2023a), we use a subset of 575 samples provide by Li (2023) for evaluation. We use all dialogues except the final round as the prompt, and the human’s question in the last round as the question.
C.2 Baseline Models
@@
(iv) GPT-4. We designed five sets of prompts for GPT-4 (Achiam et al., 2023) to inspire its ability on prompt compression and reported the best scores. Appendix B displays the prompts we employed.
C.3 Evaluation Metrics
@@
Rouge. To measure the similarity between the output of the original prompt and the compressed prompt, we apply commonly used overlap metric ROUGE (Lin, 2004). We report uni-gram and bi-gram overlap as the metric of assessing informativeness (Rouge-1 and Rouge-2), and the longest common sub-sequence as the metric of assessing fluency (Rouge-L).
@@
Appendix D Additional Results
D.1 Interpretability of Compressed prompts (Examples)
In this section, we report some examples prompts along with prompts compressed by Prompt-SAW and LLMLingua (Jiang et al., 2023a) for a quick comparison in terms of readability and/or interpretability of the compressed prompt. As an example, for the compressed prompt text compressed using LLMLingua in Table D.1, the text ”won twoes forg01 theate women prize:ertMayer” is hard to interpret for humans. On the contrary, the prompt compressed by Prompt-SAW yields comprehensive information units helpful that are not only easy to interpret but are also highly relevant to the question.
@@ D.1 Interpretability of Compressed prompts (Examples)
Document [1](Title: OPEC) Organization of the Petroleum Exporting Countries (OPEC, OH-pek, or OPEP in several other languages) is an intergovernmental organization of 14 nations as of February 2018, founded in 1960 in Baghdad by the first five members (Iran, Iraq, Kuwait, Saudi Arabia, and Venezuela), and headquartered since 1965 in Vienna, Austria. As of 2016, the 14 countries accounted for an estimated 44 percent of global oil production and 73 percent of the world’s p̈rovenöil reserves, giving OPEC a major influence on global oil prices that were previously determined by American-dominated multinational oil companies. Question: how many countries are a part of opec Answer: Compressed Prompt by LLMLingua: